# Deep Fourier-embedded Network for RGB and Thermal Salient Object Detection

### v2

@@ IV-C Ablation Study
Subsequently, we replaced FRCAB with Residual Channel Attention Block (RCAB) [76], labeled “w/ RCAB” in Table III and Fig. 8. Comparing the $7^{th}$ and last rows of Table III, we found that while the RCAB-equipped model performs well, it consistently falls short of the FRCAB-equipped model across all metrics. This clearly demonstrates that FRCAB can stably improve the model’s robustness and performance compared to RCAB in our DFENet.
@@
As shown in Fig. 9, the features fused by MPA are further refined and decoded by FRCAB. By acquiring channel-wise dependencies, FRCAB enables the model to prioritize high-frequency information and gradually generate high-resolution features.
4) Effectiveness of FEB (Sec. III-C): To validate the gains made by the proposed FEB, we replaced its core component Edge Frequency Extraction Block (EFEB) with $DNR(\cdot)$ in Eq. (21), while retaining the FRCAB in it, and presented the results in the $8^{th}$ row of Table III. The comparison revealed that without EFEB, all metrics of the model exhibit a decrease to varying degrees. Furthermore, the visualization results depicted in Fig. 8 failed to portray complete object boundaries. In contrast, as illustrated in Fig. 9(a), the complete FEB can consistently provide accurate edge features, regardless of the complexity of the scene. This ability to capture precise edge information offers a powerful boost to the model.
@@
5) Effectiveness of CFL (Sec. III-E): To evaluate the contribution of CFL, we removed it and kept only BCE to supervise the learning of the edge map, denoted as “w/o CFL” in Table III, Fig. 8 and 11. The results showed a degradation in performance on the three datasets after losing the supervision of CFL.
@@
Moreover, we displayed the 3D frequency spectrum of the results obtained with both settings in Fig. 10. It can be seen that 3D spectrograms constructed without CFL supervision exhibit oscillations and frequency distributions that deviate further from the ground-truth. Conversely, the presence of CFL optimizes the loss of frequency reconstruction and brings the frequency statistics closer to ground-truth.
@@
Furthermore, we presented a visual comparison of the edge maps for the two settings in Fig. 11. Even when both modalities provide clear objects (i.e., cases in the $1^{st}$ and $2^{nd}$ rows), the model trained only with BCE still focuses more on edge locations with high priority, making it difficult to detect complete object edges. In contrast, the model supervised with additional CFL not only demonstrates trustworthy performance in typical scenarios but also presents explicit object edges by learning difficult frequencies and reducing frequency gaps, even when object edges are difficult to distinguish from cluttered backgrounds.
Figure 12: Failure cases of the proposed DFENet. GT denotes ground-truth.
In conclusion, these findings strongly demonstrate the effectiveness of CFL in learning accurate edge maps and its importance in DFENet.
6) Increase input resolution: To verify the effectiveness of DFENet in handling high-resolution bi-modal images, we increased the input resolution to $512^{2}$. A comparison of the last two rows in Table III showed that the $512^{2}$ resolution model has comparable performance to the $384^{2}$ resolution model on the VT1000, VT5000, and VI-RGBT1500 datasets, and a slight degradation on the VT821 dataset. This performance degradation is mainly caused by the small input resolution of the pre-trained encoder used in our practice. Notably, this is the first attempt at high-resolution RGB-T SOD, whereas existing RGB-T SOD models typically have input resolution ranging from $224^{2}$ to $384^{2}$.
IV-D Complexity Analysis
@@
IV-E Limitations and Future Works
Despite achieving outstanding performance, our DFENet still struggles with some complex scenarios. As shown in Fig. 12, DFENet erroneously identifies salient objects when the background exhibits more discriminative features than foreground objects, such as higher regional contrast, greater temperature change, and faster gradient change. In contrast, due to the existence of the category concept, it is easier for the human attention mechanism to focus on central objects or living organisms while ignoring cluttered backgrounds. To mimic this mechanism, future improvements could involve adding prompts such as text and audio, or introducing large models to get more discriminating information.
Furthermore, while FFT-based operations demonstrate remarkable potential, their integration with spatial-domain operations remains an open challenge due to compatibility issues. Integrating spatial-domain techniques alongside FFT could enhance the model’s ability to capture local features and handle irregular object shapes. Therefore, combining FFT- and spatial domain-based techniques holds promising research prospects and merits further attention.
@@ V Conclusion
In this paper, we proposed DFENet to explore the potential of fast Fourier Transform (FFT) in solving RGB and thermal salient object detection (RGB-T SOD). Specifically, to fully bridge the complementary information and fuse the RGB and thermal modalities, we designed the Modal-coordinated Perception Attention (MPA), which embeds channel FFT into spatial Fourier components, achieving deep-level and multi-dimensional representation enhancement. Compared to conventional multi-head self-attention-based operations, MPA can be easily applied to high-resolution bi-modal feature fusion. In the decoding stage, we proposed the Frequency-decomposed Edge-aware Block (FEB) and Fourier Residual Channel Attention Block (FRCAB) to achieve high-resolution salient maps. Based on the analysis of bi-modal frequency decomposition, we presented FEB to clarify accurate object edges from the background and guide the integration of decoder features. We then deployed FRCAB in each decoder layer to focus on high-frequency information and obtain global dependencies in the channel dimension of the frequency domain. In the training stage, we proposed the Co-focus Frequency Loss (CFL) to assist FEB in solving hard frequencies and enhance edge representations within the bi-modal feature spaces. This frequency-domain learning synergized with spatial-domain learning to realize high-quality salient maps. Extensive experiments demonstrated that DFENet surpassed fifteen existing RGB-T SOD models on four benchmark datasets and exhaustively exhibited the representation capabilities of FFT-based operations, providing a new paradigm for a series of SOD tasks.

### v3

@@ IV-E Ablation Study
Subsequently, we replaced FRCAB with the RCAB [84], labeled “w/ RCAB” in Table III and Fig. 9. Comparing the $9^{th}$ and last rows of Table III, we found that while the RCAB-equipped model performs well, it consistently falls short of the FRCAB-equipped model across all metrics. Additionally, as shown in Fig. 10, the features processed by RCAB exhibit attention drift and edge blurring, whereas those processed by FRCAB effectively decode the fused features after MPA and focus more accurately on salient regions. By modeling channel-wise dependencies, FRCAB enables the model to prioritize high-frequency information and gradually generate high-resolution features. This clearly demonstrates that FRCAB can stably improve the model’s robustness and performance compared to RCAB in our FreqSal.
5) Effectiveness of FEB (Sec. III-C): To validate the gains made by the proposed FEB, we replaced its core component EFEB with $DNR(\cdot)$ in Eq. (21), while retaining FRCAB, and presented the results in the $10^{th}$ row of Table III. The comparison reveals that without EFEB, all metrics of the model exhibit a decrease to varying degrees. Furthermore, the visualization results depicted in Fig. 9 fail to portray complete object boundaries. In contrast, as illustrated in Fig. 10(a), the complete FEB consistently provides accurate edge features, regardless of the complexity of the scene. This ability to capture precise edge information offers a powerful boost to the model.
@@
6) Effectiveness of CFL (Sec. III-E): To evaluate the contribution of CFL, we removed it and kept only BCE to supervise edge learning, denoted as “w/o CFL” in Table III, Fig. 9 and 12. The results show a degradation in performance on the four datasets after losing the supervision of CFL.
@@
Moreover, we displayed the 3D frequency spectra of the results obtained with both settings in Fig. 11. It can be seen that 3D spectrograms constructed without CFL supervision exhibit oscillations and frequency distributions that deviate further from the ground-truth. Conversely, the presence of CFL optimizes the loss of frequency reconstruction and brings the frequency statistics closer to ground-truth.
@@
Furthermore, we presented a visual comparison of the edge maps for the two settings in Fig. 12. Even when both modalities provide clear objects (i.e., cases in the $1^{st}$-$2^{nd}$ rows), the model trained only with BCE still focuses more on edge locations with high priority, making it difficult to detect complete object edges.
TABLE IV: Quantitative Comparison of the Proposed FreqSal with High-resolution Inputs. The Best Results are Labeled Red. FPS Was Tested on an NVIDIA RTX4060 Laptop GPU.
Input Size VT821 VT1000 VT5000 VI-RGBT1500 Param(M) FLOPs(G) FPS
5122 0.869 0.024 0.903 0.015 0.885 0.023 0.891 0.026 149.6 248.0 12.6
7682 0.843 0.031 0.882 0.022 0.860 0.029 0.876 0.033 149.6 558.0 4.8
3842 0.872 0.023 0.903 0.014 0.884 0.023 0.893 0.026 149.6 139.5 17.1
In contrast, the model supervised with additional CFL not only demonstrates trustworthy performance in typical scenarios but also presents explicit object edges by learning difficult frequencies, even when object edges are difficult to distinguish from complex scenarios (i.e. $3^{rd}$-$4^{th}$ rows). In conclusion, these findings strongly demonstrate the effectiveness of CFL in learning accurate edge maps and its essential role in FreqSal.
7) Increase input resolution: To verify the effectiveness of FreqSal in handling high-resolution bimodal images, we increased the input resolution to $512^{2}$ and $768^{2}$ (trained on an NVIDIA RTX3090 GPU with 24GB of memory). A comparison in Table IV showed that the $512^{2}$ resolution model has comparable performance to the $384^{2}$ resolution model, while the $768^{2}$ resolution model experiences a noticeable performance decline. This performance degradation is likely due to the lower input resolution of the pre-trained encoder and the smaller batch size adopted during training. Significantly, this experiment represents the first exploration of high-resolution RGB-T SOD, whereas prior models typically operate within input resolution ranging from $224^{2}$ to $384^{2}$.
IV-F Generalization Experiments
1) RGB-D-T SOD: We compared our FreqSal with thirteen existing models, which consist of (1) ten bimodal models, namely CSRNet [69], CGFNet [59], DCNet [63], SwinNet [21], LSNet [40], HRTNet [22], PICRNet [57], LAFB [62], MAGNet [72], and UniTR [28], (2) three trimodal models, namely HWSI [86], MFFNet [96], and QSFNet [97]. To ensure a fair comparison, all saliency maps used for comparison were obtained from the corresponding models’ homepages or generated using open-source code.
TABLE V: Quantitative Comparison of Our FreqSal and Thirteen State-of-the-Art Models on the VDT-2048 dataset. VT, VD, and VDT Denote Training and Testing on RGB-T, RGB-D, and RGB-D-T Data, Respectively. The Best Results are Labeled Red and the Second Best Results are Labeled Blue.
Model Tpye Backbone VDT2048
$Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$
Bimodal Model
CSRNet21 [69] VT ESPNetv2 0.9494 0.8821 0.8099 0.7888 0.0050
CGFNet21 [59] VT VGG-16 0.9319 0.9166 0.8425 0.7822 0.0033
DCNet22 [63] VT VGG-16 0.9658 0.8787 0.8227 0.8521 0.0038
SwinNet22 [21] VT Swin-B 0.9444 0.9370 0.8841 0.8090 0.0026
SwinNet22 [21] VD Swin-B 0.8978 0.9198 0.8344 0.7321 0.0037
LSNet23 [40] VT MobileNetv2 0.9327 0.8867 0.8005 0.7607 0.0044
HRTNet23 [22] VT HRFormer 0.9680 0.9281 0.8840 0.8446 0.0026
HRTNet23 [22] VD HRFormer 0.9617 0.9144 0.8603 0.8227 0.0031
PICRNet23 [57] VD Swin-T 0.9574 0.8941 0.8267 0.7998 0.0040
LAFB24 [62] VT Res2Net-50 0.9736 0.9325 0.8943 0.8554 0.0025
LAFB24 [62] VD Res2Net-50 0.9676 0.9160 0.8668 0.8356 0.0031
MAGNet24 [72] VT SMT 0.9758 0.9266 0.8848 0.8563 0.0028
MAGNet24 [72] VD SMT 0.9647 0.9092 0.8533 0.8222 0.0034
UniTR24 [28] VT Swin-B 0.9805 0.9084 0.8704 0.8789 0.0028
UniTR24 [28] VD Swin-B 0.9767 0.8930 0.8420 0.8546 0.0034
FreqSal (Ours) VT CDFFormer-M 0.9862 0.9384 0.8954 0.8809 0.0024
Trimodal Model
HWSI22 [86] VDT VGG-16 0.9815 0.9318 0.8974 0.8718 0.0026
MFFNet24 [96] VDT VGG-16 0.9831 0.9394 0.9017 0.8758 0.0025
QSFNet24 [97] VDT Swin-B 0.9868 0.9426 0.9145 0.8902 0.0021
FreqSal (Ours) VT CDFFormer-M 0.9862 0.9384 0.8954 0.8809 0.0024
Table V and Fig. 13 present the quantitative comparison of the proposed FreqSal with thirteen SOTA models on the VDT-2048 dataset. In summary, our FreqSal surpassed ten existing bimodal models across all six metrics and achieved performance comparable to trimodal models while using only two modalities. Specifically, compared to the leading bimodal models on RGB-T data, FreqSal surpassed UniTR [28] by 0.6% and 0.2% in $Em$ and $F_{\beta}$, respectively, and exceeded LAFB [62] by 0.1% and 4.0% in $wF_{\beta}$ and $\mathcal{M}$, respectively. Compared to trimodal models, FreqSal trained on RGB-T data solidly outperformed HWSI [86] and led MFFNet [96] in $Em$, $F_{\beta}$, and $\mathcal{M}$.
Figure 13: (a) F-measure-threshold (FT) and (b) Precision-recall (PR) curves of our FreqSal and existing ten bimodal SOD models on the VDT-2048 dataset.
Although it lagged behind QSFNet [97], our FreqSal remained competitive with trimodal models. These findings highlight the effectiveness and strong generalization ability of our FreqSal on the VDT-2048 dataset.
2) RGB-D SOD: We conducted a comprehensive comparison with sixteen existing RGB-D SOD models, namely LSNet [40], PopNet [98], C2DFNet [99], HiDANet [47], RD3D+ [100], FasterSal [101], LAFB [62], VST [6], SwinNet [21], CAVER [24], PICRNet [57], MAGNet [72], HFMDNet [56], EM-Trans [102], VST++ [27], and UniTR [28]. To ensure a fair comparison, all saliency maps were either obtained from the authors of the corresponding models or generated using their publicly available code.
Table VI and Fig. 14 present the comparison of the proposed FreqSal with sixteen RGB-D SOD models on the NLPR, NJUD, DUT-RGBD, SIP, and STERE datasets using six evaluation metrics: $Em$, SmSm, $wF_{\beta}$, $F_{\beta}$, $\mathcal{M}$, and PR curve. Overall, our FreqSal surpassed sixteen existing RGB-D models across all six evaluation metrics on the five datasets. Specifically, FreqSal outperformed the pervious best-performing model, VST++ [27], achieving average improvements of 0.9%, 0.4%, 1.1%, 0.5%, 7.0% in $Em$, SmSm, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$, respectively. Compared to HFMDNet [56], FreqSal shows a substantial performance advantage on the SIP and STERE datasets, while maintaining comparable results on the NLPR and NJUD datasets. This strong performance highlights the generalization ability of FreqSal’s deep Fourier-embedded strategy, which proves equally effective in addressing the RGB-D SOD task.
TABLE VI: Quantitative Comparison of Our FreqSal and Sixteen State-of-the-Art RGB-D SOD Models on the NLPR, NJUD, DUT-RGBD, SIP, and STERE datasets. The Best Results are Labeled Red and the Second Best Results are Labeled Blue.
Model NLPR NJUD DUT-RGBD SIP STERE
$Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$
CNN-based Model
LSNet23 [40] 0.955 0.919 0.873 0.881 0.025 0.922 0.911 0.879 0.899 0.038 - - - - - 0.927 0.886 0.848 0.882 0.049 0.913 0.871 0.817 0.854 0.054
PopNet23 [98] 0.961 0.936 0.904 0.903 0.020 0.929 0.925 0.904 0.916 0.031 - - - - - 0.929 0.894 0.867 0.890 0.043 0.931 0.920 0.889 0.903 0.033
C2DFNet23 [99] 0.958 0.928 0.890 0.897 0.021 0.919 0.908 0.879 0.898 0.038 0.958 0.933 0.919 0.933 0.025 0.916 0.871 0.831 0.865 0.053 0.926 0.902 0.863 0.881 0.038
HiDANet23 [47] 0.959 0.930 0.898 0.899 0.022 0.935 0.926 0.908 0.919 0.030 - - - - - 0.925 0.893 0.864 0.889 0.044 0.934 0.912 0.880 0.894 0.036
RD3D+24 [100] 0.958 0.933 0.883 0.886 0.022 0.920 0.928 0.894 0.910 0.033 0.953 0.936 0.909 0.945 0.031 0.924 0.892 0.850 0.883 0.046 0.921 0.914 0.860 0.881 0.039
FasterSal24 [101] 0.955 0.922 0.890 0.895 0.023 0.919 0.909 0.891 0.902 0.034 0.952 0.920 0.909 0.922 0.031 0.924 0.872 0.841 0.867 0.049 0.918 0.890 0.856 0.872 0.040
LAFB24 [62] 0.958 0.930 0.895 0.902 0.022 0.924 0.924 0.904 0.918 0.029 0.957 0.931 0.915 0.929 0.027 0.937 0.897 0.875 0.904 0.041 0.930 0.907 0.875 0.895 0.037
Transformer-based Model
VST21 [6] 0.954 0.931 0.887 0.886 0.023 0.914 0.922 0.888 0.900 0.034 0.960 0.943 0.920 0.921 0.025 0.937 0.904 0.873 0.895 0.040 0.917 0.913 0.866 0.878 0.038
SwinNet22 [21] 0.967 0.941 0.908 0.908 0.018 0.934 0.935 0.913 0.922 0.027 0.968 0.949 0.935 0.943 0.020 0.943 0.911 0.890 0.912 0.035 0.929 0.919 0.882 0.893 0.033
CAVER23 [24] 0.964 0.934 0.904 0.906 0.021 0.931 0.927 0.909 0.920 0.028 0.962 0.937 0.926 0.936 0.026 0.937 0.904 0.879 0.898 0.037 0.935 0.917 0.888 0.901 0.032
PICRNet23 [57] 0.965 0.935 0.904 0.905 0.019 0.932 0.927 0.907 0.917 0.029 0.965 0.944 0.933 0.944 0.021 0.933 0.899 0.875 0.897 0.040 0.936 0.920 0.891 0.903 0.031
MAGNet24 [72] 0.964 0.938 0.910 0.912 0.018 0.928 0.929 0.912 0.922 0.027 0.966 0.944 0.935 0.946 0.021 0.941 0.907 0.890 0.911 0.036 0.930 0.922 0.893 0.903 0.030
HFMDNet24 [56] 0.968 0.938 0.916 0.924 0.017 0.939 0.937 0.928 0.937 0.023 - - - - - 0.925 0.886 0.868 0.897 0.044 0.933 0.918 0.892 0.905 0.031
EM-Trans24 [102] 0.966 0.939 0.913 0.919 0.017 0.934 0.930 0.915 0.926 0.027 - - - - - 0.938 0.904 0.886 0.912 0.039 0.938 0.925 0.900 0.913 0.028
VST++24 [27] 0.963 0.938 0.910 0.912 0.020 0.941 0.940 0.925 0.933 0.024 0.969 0.953 0.942 0.943 0.020 0.942 0.912 0.893 0.911 0.034 0.928 0.926 0.896 0.906 0.030
UniTR24 [28] 0.970 0.930 0.912 0.925 0.018 0.940 0.925 0.918 0.931 0.025 - - - - - 0.940 0.896 0.888 0.915 0.037 0.933 0.907 0.884 0.897 0.032
FreqSal (Ours) 0.974 0.944 0.922 0.932 0.017 0.944 0.943 0.930 0.912 0.023 0.973 0.952 0.946 0.948 0.019 0.950 0.917 0.907 0.927 0.033 0.945 0.932 0.909 0.909 0.027
Figure 14: F-measure curves and precision-recall (PR) curves of the proposed FreqSal and sixteen existing RGB-D SOD models on the NLPR, NJUD, DUT-RGBD, SIP, and STERE datasets.
IV-G Complexity Analysis
@@
IV-H Limitations and Future Works
Despite achieving outstanding performance, our FreqSal still struggles with some complex scenarios.
Figure 15: Failure cases of the proposed FreqSal. GT denotes ground-truth.
As shown in Fig. 15, FreqSal erroneously identifies salient objects when the background exhibits more discriminative features than foreground objects, such as higher regional contrast and greater temperature change. In contrast, due to the existence of the category concept, it is easier for the human attention mechanism to focus on central objects or living organisms while ignoring cluttered backgrounds. To mimic this mechanism, future improvements could involve adding prompts such as text and audio, or introducing large models to get more discriminating information about salient objects.
Furthermore, while FFT-based operations demonstrate remarkable potential, their integration with spatial-domain operations remains an open challenge due to compatibility issues. The combination of frequency- and spatial-domain techniques improves the model’s ability to capture irregular object shapes, highlighting strong potential for future research. This direction can be further advanced by designing normalization layers and optimizers tailored to frequency-domain features as well as by establishing a comprehensive theoretical framework.
@@ V Conclusion
In this paper, we proposed FreqSal to explore the potential of FFT in solving RGB-T SOD. Specifically, to fully bridge the complementary information and fuse the RGB and thermal modalities, we designed the MPA, which embeds channel FFT into spatial Fourier components, achieving deep-level and multi-dimensional representation enhancement. Compared to conventional Transformer-based operations, MPA can be easily applied to high-resolution bimodal feature fusion. In the decoding stage, we proposed the FEB and FRCAB to achieve high-resolution saliency maps. Based on the analysis of bimodal frequency decomposition, we presented FEB to clarify accurate object edges from the background and guide the integration of decoder features. We then deployed FRCAB in each decoder layer to focus on high-frequency information and obtain global dependencies in the channel dimension of the frequency domain. In the training stage, we proposed the CFL to assist FEB in solving hard frequencies and enhance edge representations within the bimodal feature spaces. This frequency-domain learning synergized with spatial-domain learning to realize high-quality saliency maps. Extensive experiments demonstrate that FreqSal not only surpasses twenty-nine existing bimodal models on the ten publicly available benchmark datasets but more importantly, reveals the under-explored potential of Fourier-domain representations, potentially paving the way for paradigm evolution in the SOD tasks.
Appendix
V-A Summary of Transformer-based Accuracy-targeted Model Structures
We summarized the simplified architectures of existing Transformer-based accuracy-targeted RGB-T SOD models in Fig. 16. To reduce memory usage when training under limited computational resources, these models typically adopt three compromise strategies: (1) VST [6], PRLNet [25], VST++ [27], UniTR [28], and SACNet [29] ignore high-resolution bimodal features; (2) IFFNet [20], SwinNet [21], CGMDRNet [23], HRTNet [22], and CACFNet [30] employ CNN-based fusion modules; and (3) CGMDRNet [23], CACFNet [30], and CAVER [24] employ CNN-based pretrained encoders. Unlike these Transformer-based models, our FreqSal not only effectively fuses high-resolution bimodal features but also employs FFT-based designs to capture global relationships across all stages.
Figure 16: Simplified structure diagrams of existing Transformer-based accuracy-targeted RGB-T SOD models. (a) represents models that ignore high-resolution bimodal features, including VST [6], PRLNet [25], VST++ [27], UniTR [28], and SACNet [29]. (b) presents models that employ CNN-based fusion modules for bimodal feature interaction, including IFFNet [20], SwinNet [21], CGMDRNet [23], HRTNet [22], and CACFNet [30]. (c) shows models that employ CNN-based pretrained encoders to extract initial bimodal features, including CGMDRNet [23], CACFNet [30], and CAVER [24]. Unlike these current models, all components of our FreqSal are designed using FFT (d), allowing it to capture global relationships across all stages.
V-B Additional Ablation Study
1) Effectiveness of activation function in MPA (Sec. III-B): To validate the effectiveness of StarReLU [83] as the activation function in the proposed MPA, we compared it with two standard activations, namely ReLU and GeLU, labeled “StarReLU $\rightarrow$ ReLU” and “StarReLU $\rightarrow$ GeLU” in Table VII. The variants with ReLU and GeLU performed worse across all four datasets, confirming StarReLU’s superiority. Compared with ReLU and GeLU, StarReLU produces smoother frequency-domain responses and more stable gradient behavior, as its squaring operation corresponds to the square of the amplitude in the Fourier domain, thereby preventing abrupt truncation of high-frequency components and preserving phase information that mainly resides in these frequencies. Moreover, the inclusion of a bias term allows weak activations in negative regions, further supporting phase preservation. In addition, StarReLU grows linearly in its non-zero regions, effectively avoiding gradient explosion—an essential property for stable inverse Fourier transforms and gradient backpropagation in the frequency domain. Therefore, StarReLU proves more suitable for the proposed MPA, which involves successive Fourier transforms.
TABLE VII: Ablation Study of Activation Function in Modal-coordinated Perception Attention. The Best Results are Labeled Red.
Settings VT821 VT1000 VT5000 VI-RGBT1500
StarReLU $\rightarrow$ ReLU 0.868 0.024 0.897 0.015 0.879 0.024 0.886 0.027
StarReLU $\rightarrow$ GeLU 0.863 0.023 0.893 0.014 0.881 0.024 0.884 0.026
StarReLU 0.872 0.023 0.903 0.014 0.884 0.023 0.893 0.026
2) Effectiveness of the input of FEB (Sec. III-C): To validate the effectiveness of the input selection strategy for FEB, we compared the original setting (i.e., “T” and “F” as inputs) with three representative settings, and presented the results in Table VIII. Where “RGB”, “T”, and “F” denote RGB, thermal, and fusion features, respectively. The results show that the best performance is achieved when the fusion and thermal features serve as inputs. In contrast, incorporating RGB features leads to a significant drop in $F_{\beta}$ metrics. As illustrated in Fig. 4, the phase and high-frequency components of RGB images contain not only edge information but also predominantly cluttered background textures. Since edge and texture frequency signals coexist within the same frequency band, introducing RGB features as FEB inputs dilutes edge signals with background noise and leads to blurred edge features or edge misdetection. In summary, we selected thermal and fusion features as inputs for the FEB to obtain clear edge features.
TABLE VIII: Ablation Study of the Inputs of Frequency-decomposed Edge-aware Block. The Best Results are Labeled Red.
Setting VT821 VT1000 VT5000 VI-RGBT1500
T 0.870 0.024 0.899 0.015 0.878 0.024 0.892 0.027
F 0.865 0.025 0.892 0.016 0.876 0.024 0.890 0.027
RGB + T + F 0.858 0.023 0.882 0.017 0.871 0.024 0.886 0.026
T + F 0.872 0.023 0.903 0.014 0.884 0.023 0.893 0.026
3) Selection of weight parameters $\left\{{{\lambda_{i}}}\right\}_{i=1}^{4}$ in Eq. (42) (Sec. III-E): For the weight parameters $\left\{{{\lambda_{i}}}\right\}_{i=1}^{4}$ of loss functions in Eq. (42), we determined their values through careful testing and set them to 1 in this paper. The representative values and corresponding results are presented in Table IX. Specifically, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ respectively control the importance of the saliency loss ${\cal L}_{S}$}, the decoding loss ${\cal L}_{D}$}, the spatial-domain edge loss ${\cal L}_{E}$}, and the co-focus loss ${\cal L}_{CFL}$} in the total loss ${\cal L}_{total}$}. When the values of $\lambda_{1}$ or $\lambda_{2}$ deviate from 1, the model performance generally declines to varying degrees. For edge feature learning in the bi-domain, a higher spatial-domain loss proportion (i.e., $\lambda_{3}$ is greater than $\lambda_{4}$) improves the $\mathcal{M}$ metric, while a higher frequency-domain loss proportion (i.e., $\lambda_{4}$ is greater than $\lambda_{3}$) enhances the $F_{\beta}$ metric. When all $\left\{{{\lambda_{i}}}\right\}_{i=1}^{4}$ are set to 1, the model achieves the most balanced performance across all four datasets.
TABLE IX: Selection of Weight Parameters $\left\{{{\lambda_{i}}}\right\}_{i=1}^{4}$ in Eq. (42). The Best Results are Labeled Red.
Parameters VT821 VT1000 VT5000 VI-RGBT1500
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$
0.5 1 1 1 0.868 0.025 0.897 0.015 0.877 0.024 0.890 0.027
1.5 1 1 1 0.870 0.024 0.905 0.015 0.882 0.024 0.893 0.026
1 0.5 1 1 0.870 0.024 0.899 0.014 0.882 0.023 0.890 0.026
1 1.5 1 1 0.872 0.024 0.906 0.015 0.886 0.024 0.889 0.027
1 1 0.5 1 0.876 0.024 0.906 0.015 0.886 0.024 0.895 0.027
1 1 1.5 1 0.868 0.024 0.903 0.014 0.880 0.023 0.889 0.026
1 1 1 0.5 0.867 0.023 0.905 0.015 0.878 0.024 0.889 0.027
1 1 1 1.5 0.874 0.024 0.899 0.016 0.887 0.024 0.896 0.028
1 1 1 1 0.872 0.023 0.903 0.014 0.884 0.023 0.893 0.026
V-C Additional Discussion
In the paper, we primarily conducted experiments on FreqSal to evaluate core metrics (e.g., $Em$, SmSm, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$) for the bimodal SOD task, which, while indirectly reflecting a model’s use of global context, do not explicitly assess its spatial reasoning ability.
In future work, we plan to develop the FFT-based approach and evaluate it on benchmarks specifically designed for spatial reasoning. By examining the model’s ability to understand spatial relationships, including relative object positions and scene-level structural patterns, we aim to provide direct quantitative evidence supporting the generalizability of FFT-based methods in capturing global dependencies.